A business executive has the option of investing money in two plans. Plan A guarantees that each dollar invested will earn 70 cents a year hence, and plan B guarantees that each dollar invested will earn $2 two years hence. Plan A allows yearly

investments, while in plan B, only investments for periods that are multiples of two

years are allowed. How should the executive invest $100,000 to maximize the earnings at the end of three years? Formulate this problem as a linear programming

problem


1
Expert's answer
2021-11-17T12:25:09-0500

Let XIA=Amount invested in year i using plan A,i=1,2,3


Let XIB=Amount invested in year i using plan B,i=1,2,3


Let us first formulate the constraint as follows: 

X1A+X 1B"=" 100,000

The earning at the end of the first year is X1A+0.7X1A=1.7X1A


Investment at the beginning of the 2nd year is X2A+X2B ,which must be equal to the earning at the end of the first year, that is;

X2A+X2B=1.7X1A


The earning at the end of the 2nd year is 3X1B +1.7X2A


Investment at the beginning of the 3rd year is X3A ,which must be equal to the earning at the end of the 2nd year, that is, 3X1B +1.7X2A=X3A


The total earning at the end of the 3rd year is 3x2B+1.7x3A


Therefore, LP model of the problem is ;


Maximize Z=3x2B+1.7x3A


Subject to;

X1A+X 1B"=" 100,000


X2A+X2B=1.7X1A


3X1B +1.7X2A=X3A


X1A , X1B, X2A,X2B, X3A"\ge" 0